# Robust and Computation-Aware Gaussian Processes

### v1

@@ 1 Introduction
Gaussian Processes (GPs) are a foundational tool in probabilistic machine learning, offering non-parametric modeling with principled uncertainty estimates [35]. Their use spans diverse domains such as time series forecasting, spatial modeling, and regression on structured scientific data. One particularly impactful area where GPs have become integral is Bayesian Optimization (BO), a sample-efficient framework for optimizing expensive black-box functions, where the surrogate model’s predictive accuracy and uncertainty calibration directly influence decision-making [10]. Many scientific and industrial tasks are high-throughput optimization problems in which numerous experiments or simulations are performed, possibly in parallel. These can be effectively tackled by BO and include material design [14], drug discovery with extensive chemical space search [17], and fine-tuning gene expression in synthetic biology [5].
Yet, two major challenges limit the reliability of GPs in both regression and BO settings: scalability to large datasets and robustness to corrupted or outlier-contaminated observations. Sparse Variational Gaussian Processes (SVGPs, [16, 40, 41]) address scalability via inducing point approximations, yielding a low-rank approximation that allows large-scale regression and high-throughput BO.
@@
Despite their computational efficiency, significant challenges hinder the adoption of SVGPs for high-throughput BO tasks, due to the impact of surrogate posterior quality on data acquisition. For starters, SVGP is not robust against outliers: upon contamination of the observed data, the posterior mean can deviate significantly from the true latent function. Such deviation will misguide most acquisition strategies, resulting in suboptimal performance. Outliers are an almost unavoidable aspect of many real-world applications and datasets, specifically in high-throughput settings. They can arise from various factors, such as faulty measurements, malfunctioning sensors, or genetic anomalies. Next, SVGPs suffer from overconfident predictions, a well-known pathology caused by the error incurred by the posterior approximation. Overconfident predictions occur in regions where the inducing points are far from observed data. Accurate uncertainty quantification is essential for sequentially identifying the optimal data points in BO [44].
Contributions. To address these challenges, we propose a novel approach that combines robust-conjugate SVGP [2] with computation-aware GP [46]. This method, referred to as the Robust Computation-Aware GP (RCaGP), jointly tackles outlier robustness and approximation-induced uncertainty—two challenges that are deeply intertwined in sparse GP models. Computational approximations such as low-rank matrix factorizations can exacerbate the effects of outliers, and mitigating one without addressing the other leads to suboptimal performance. RCaGP can be used as a surrogate in high-throughput BO, and its design is motivated by the need to address these two problems in tandem. This can be seen in Figure 1. In summary, this work
@@
    (Methodological) introduces a novel GP model (RCaGP) that jointly addresses approximation-awareness and outlier robustness through a principled probabilistic framework—going beyond previous work that tackles these issues in isolation.
@@ 2 Background
Robust conjugate GP. The Gaussian noise assumption in GPs often leads to misspecification, e.g., due to outliers. To address this problem, recent work proposed robust conjugate GP (RCGP), which employs a robust loss function through generalized Bayesian inference [2]. Consider a noisy dataset ${\mathcal{D}}_{n}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}=({\bf{X}},{\bf{y}})$. We specify a GP prior for the finite function evaluations ${\bf{f}}=[f({\bf{x}}_{1}),\dots,f({\bf{x}}_{n})]$: $p({\bf{f}})={\mathcal{N}}({\bf{f}};\!{\bf{m}},\!{\bf{K}})$, with ${\bf{m}}=[m({\bf{x}}_{1}),\dots,m({\bf{x}}_{n})]^{\top}$ and ${\bf{K}}=[k({\bf{x}}_{j},\!{\bf{x}}_{l})]_{1\leq j,l\leq n}$. Both $m:{\mathcal{X}}\rightarrow$ and $k:{\mathcal{X}}\times{\mathcal{X}}\rightarrow$ denote the mean and kernel functions, encoding prior beliefs about the function’s structure and smoothness [35]. Conditioning on observations, the posterior predictive distribution is Gaussian with mean $\mu_{\ast}({\bf{x}})$ and variance $k_{\ast}({\bf{x}},\!{\bf{x}})$.
@@
Nevertheless, RCGP inherits the cubic inference complexity of vanilla GPs, prohibiting its usage on large-scale datasets like those encountered in high-throughput BO. To remedy this issue, RCSVGP were introduced, enabling outlier-robust regression on large-scale data [2]. This model introduces inducing points ${\bf{u}}=\{{\bf{u}}_{l}\}_{l=1}^{r}$ for $r\ll n$, with inducing locations ${\bf{z}}=\{{\bf{z}}_{l}\}_{l=1}^{r}$ such that ${\bf{u}}_{l}=f({\bf{z}}_{l})$. Given ${\bf{u}}$, RCSVGP constructs an approximate posterior $q({\bf{u}})={\mathcal{N}}({\boldsymbol{\mu}}_{\bf{u}},{\boldsymbol{\Sigma}}_{\bf{u}})$ with variational parameters ${\boldsymbol{\mu}}_{\bf{u}}$ and ${\boldsymbol{\Sigma}}_{\bf{u}}$, obtained by maximizing the evidence lower bound (ELBO) criterion:
@@
Computation-aware GP. [45, 46] proposed computation-aware GP (CaGP) to address the SVGP overconfidence. CaGP mitigates this issue by ensuring its posterior variance exceeds the vanilla GP variance. This corrected variance is obtained by adding computational uncertainty, associated with the approximation error in the representer weights ${\bf{v}}_{\ast}=({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})^{-1}({\bf{y}}-{\bf{m}})$. To obtain computational uncertainty, CaGP places a prior $p({\bf{v}}_{\ast})={\mathcal{N}}({\bf{v}}_{0},{\boldsymbol{\Sigma}}_{0})$ and updates its posterior distribution $\mathcal{N}({\bf{v}}_{i},{\boldsymbol{\Sigma}}_{i})$ using a probabilistic linear solver based on linear Gaussian identities:
@@
where ${\bf{C}}_{i}={\bf{S}}_{i}({\bf{S}}_{i}^{\top}\hat{{\bf{K}}}^{-1}{\bf{S}}_{i}){\bf{S}}_{i}\approx{\bf{K}}^{-1}$ is a low-rank approximation and ${\bf{S}}_{i}\in^{n\times i}$ is the action matrix associated with the chosen approximation technique. In the case of inducing points, ${\bf{S}}_{i}=k({\bf{X}},{\bf{z}}_{i})$. Given a new data point ${\bf{x}}$, the predictive distribution of $f({\bf{x}})$ is Gaussian, characterized by the mean $\mu_{i}({\bf{x}})$ and variance $k_{i}({\bf{x}},{\bf{x}})$, given by:
@@
Bayesian optimization. BO is a sequential algorithm designed to solve the problem ${\mathrm{argmax}}_{{\bf{x}}\in{\mathcal{X}}}f({\bf{x}})$ for ${\mathcal{X}}$ a compact domain, and $f$ a black-box function, meaning only noisy evaluations are accessible [10]. BO involves a probabilistic surrogate $p(f|{\mathcal{D}})$ and an acquisition function (AF) $\alpha({\bf{x}}):=\int u({\bf{x}};f)\,p(f|{\mathcal{D}})\,\mathrm{d}f$, where $u$ measures the usefulness of evaluating $f({\bf{x}})$. GPs are usually employed as the surrogate, and popular AFs include, for instance, expected improvement, knowledge gradient, and entropy search [10]. Importantly, these AFs can be maximized in a batch setting, yielding $q$ candidates every iteration $t,{\bf{X}}_{t}^{q}={\mathrm{argmax}}_{{\bf{X}}\in{\mathcal{X}}^{q}}\alpha({\bf{X}})$ [48].
@@ 3.1 Robust Computation-aware Inference
Inspired by CaGP, we apply a probabilistic treatment to the representer weights $\hat{{\bf{v}}}=({\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}\!-\!{\bf{m}}_{\bf{w}})$ of RCGP in Equation 1. This enables us to quantify uncertainty in approximating $\hat{{\bf{v}}}$—an aspect most approximate GPs ignore. RCaGP later incorporates this uncertainty into its predictive variance. We begin by placing a prior on $\hat{{\bf{v}}}$, i.e., $\hat{{\bf{v}}}\sim{\mathcal{N}}(\tilde{{\bf{v}}}_{0}\!=\!\mathbf{0},\,\tilde{{\boldsymbol{\Sigma}}}_{0}\!=\!\tilde{{\bf{K}}}^{-1})$, where $\tilde{{\bf{K}}}={\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}}$. We then update this belief via the linear Gaussian identity [45], yielding a Gaussian posterior with mean $\tilde{{\bf{v}}}_{i}$ and variance $\tilde{{\boldsymbol{\Sigma}}}_{i}$:
@@
where $\tilde{{\bf{C}}}_{i}\!=\!{\bf{S}}_{i}({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}$ and ${\bf{S}}_{i}\!\in\!^{n\times i}$ are low-rank approximations of $\tilde{{\bf{K}}}^{-1}$ and the actions, respectively. The updates can be done sequentially [45, Algorithm 1] or in batches [46, Algorithm S2].
Recent work generalized the action matrix ${\bf{S}}_{i}\in^{n\times i}$ by treating it as a variable that can be optimized during model selection [46]. Given a new data point ${\bf{x}}\in{\mathcal{X}}$, the predictive distribution of $f({\bf{x}})$ follows a normal distribution with mean $\hat{\mu}_{i}({\bf{x}})$ and variance $\hat{k}_{i}({\bf{x}},{\bf{x}})$:
@@
In Equation 11, RCaGP incorporates the computational uncertainty ${\bf{k}}_{\bf{x}}^{\top}\,\tilde{{\boldsymbol{\Sigma}}}_{i}\,{\bf{k}}_{\bf{x}}$, corresponding to the variance of the representer weights $\tilde{{\boldsymbol{\Sigma}}}_{i}$, resulting in a combined uncertainty $\hat{k}_{i}({\bf{x}},{\bf{x}})$. This computational uncertainty guarantees that the predictive posterior variance of RCaGP is larger than RCGP, preventing RCaGP from overconfident predictions. Such behavior is referred to as a conservative uncertainty estimate. The computational complexity of RCaGP can be found in Appendix B.
The robustness of RCaGP hinges on the weight function $w$, which we choose to incorporate into the representer weights $\tilde{{\bf{v}}}$. To that end, we follow [2] and let:
@@
The weight function assigns smaller values to observations $y$ that deviate significantly from the mean prior $m({\bf{x}})$, treating them as potential outliers. Lower weights reduce the influence of such points during inference, leading to a more robust posterior. Additionally, the weight enters the noise term ${\bf{J}}_{\bf{w}}=\mathrm{diag}([{\sigma_{\mathrm{noise}}^{2}}/2{\bf{w}}^{-2}])$, effectively inflating the noise for outliers and further limiting their impact. $w({\bf{x}},y)$ depends critically on the mean prior: for an outlier $\hat{y}$, if $m(\hat{{\bf{x}}})$ is close to $\hat{y}$, the weight increases. However, choosing an informative mean prior is difficult without domain knowledge.
@@ 3.2 Model hyperparameters optimization
We use the evidence lower bound (ELBO) as a loss function to optimize the kernel hyperparameters ${\boldsymbol{\theta}}\!\in\!^{p}$. Following [46], this enables RCaGP to scale to large datasets while avoiding overconfidence. The variational family is defined using the RCaGP posterior $q_{i}({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})$, i.e., $\{q_{i}({\bf{f}})={\mathcal{N}}(\hat{\mu}_{i}({\bf{X}}),\hat{k}_{i}({\bf{X}},\!{\bf{X}}))|{\bf{S}}_{i}\!\in\!^{n\times i}\}$, parameterized by action ${\bf{S}}_{i}$. We replace the evidence term $\log p({\bf{y}}|{\boldsymbol{\theta}})$ with the robust loss $\log\Psi^{w}({\bf{y}},\!{\bf{f}})$ from [2]. We then formulate ELBO as
@@
This loss learns the hyperparameters as if maximizing ${\mathbb{E}}_{p({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ while minimizing the computational uncertainty associated with the approximation error. The derivation of the expected loss ${\mathbb{E}}_{q({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ follows [2], while the derivation of the KL term ${\mathrm{KL}}[q({\bf{f}})\|p({\bf{f}})]$ draws an analogy to CaGP. For the closed form and detailed derivation of ELBO, we refer the readers to Proposition C.1. We obtain the optimal hyperparameters ${\boldsymbol{\theta}}^{\ast}$ by minimizing the negative ELBO.

### v2

@@ 1 Introduction
Gaussian Processes (GPs) are a foundational tool in probabilistic machine learning, offering non-parametric modeling with principled uncertainty estimates [Rasmussen2006]. Their use spans diverse domains such as time series forecasting, spatial modeling, and regression on structured scientific data. One particularly impactful area where GPs have become integral is Bayesian Optimization (BO), a sample-efficient framework for optimizing expensive black-box functions, where the surrogate model’s predictive accuracy and uncertainty calibration directly influence decision-making [garnett2023bayesian]. Many scientific and industrial tasks are high-throughput optimization problems in which numerous experiments or simulations are performed, possibly in parallel. These can be effectively tackled by BO and include material design [hastings2024interoperable], drug discovery with extensive chemical space search [hernandez2016distributed], and fine-tuning gene expression in synthetic biology [belanger2019biological].
Yet, two major challenges limit the reliability of GPs in both regression and BO settings: scalability to large datasets and robustness to corrupted or outlier-contaminated observations. Sparse Variational Gaussian Processes (SVGPs, [hensman2013gaussian, titsias2009variational, titsias2014doubly]) address scalability via inducing point approximations, yielding a low-rank approximation that allows large-scale regression and high-throughput BO.
@@
Despite their computational efficiency, significant challenges hinder the adoption of SVGPs for high-throughput BO tasks, due to the impact of surrogate posterior quality on data acquisition. For starters, SVGP is not robust against outliers: upon contamination of the observed data, the posterior mean can deviate significantly from the true latent function. Such deviation will misguide most acquisition strategies, resulting in suboptimal performance. Outliers are an almost unavoidable aspect of many real-world applications and datasets, specifically in high-throughput settings. They can arise from various factors, such as faulty measurements, malfunctioning sensors, or genetic anomalies. Next, SVGPs suffer from overconfident predictions, a well-known pathology caused by the error incurred by the posterior approximation. Overconfident predictions occur in regions where the inducing points are far from observed data. Accurate uncertainty quantification is essential for sequentially identifying the optimal data points in BO [wang2018batched].
Contributions. To address these challenges, we propose a novel approach that combines robust-conjugate SVGP [altamirano2023robust] with computation-aware GP [wenger2024computation]. This method, referred to as the Robust Computation-Aware GP (RCaGP), jointly tackles outlier robustness and approximation-induced uncertainty, two challenges that are deeply intertwined in sparse GP models. Computational approximations such as low-rank matrix factorizations can exacerbate the effects of outliers, and mitigating one without addressing the other leads to suboptimal performance. RCaGP can be used as a surrogate in high-throughput BO, and its design is motivated by the need to address these two problems in tandem. This can be seen in Figure 1. In summary, this work
@@
    (Methodological) introduces a novel GP model (RCaGP) that jointly addresses approximation-awareness and outlier robustness through a principled probabilistic framework, going beyond previous work that tackles these issues in isolation.
@@ 2 Background
Robust conjugate GP. The Gaussian noise assumption in GPs often leads to misspecification, e.g., due to outliers. To address this problem, recent work proposed robust conjugate GP (RCGP), which employs a robust loss function through generalized Bayesian inference [altamirano2023robust]. Consider a noisy dataset ${\mathcal{D}}_{n}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}=({\bf{X}},{\bf{y}})$. We specify a GP prior for the finite function evaluations ${\bf{f}}=[f({\bf{x}}_{1}),\dots,f({\bf{x}}_{n})]$: $p({\bf{f}})={\mathcal{N}}({\bf{f}};\!{\bf{m}},\!{\bf{K}})$, with ${\bf{m}}=[m({\bf{x}}_{1}),\dots,m({\bf{x}}_{n})]^{\top}$ and ${\bf{K}}=[k({\bf{x}}_{j},\!{\bf{x}}_{l})]_{1\leq j,l\leq n}$. Both $m:{\mathcal{X}}\rightarrow$ and $k:{\mathcal{X}}\times{\mathcal{X}}\rightarrow$ denote the mean and kernel functions, encoding prior beliefs about the function’s structure and smoothness [Rasmussen2006]. Conditioning on observations, the posterior predictive distribution is Gaussian with mean $\mu_{\ast}({\bf{x}})$ and variance $k_{\ast}({\bf{x}},\!{\bf{x}})$.
@@
Nevertheless, RCGP inherits the cubic inference complexity of vanilla GPs, prohibiting its usage on large-scale datasets like those encountered in high-throughput BO. To remedy this issue, RCSVGP were introduced, enabling outlier-robust regression on large-scale data [altamirano2023robust]. This model introduces inducing points ${\bf{u}}=\{{\bf{u}}_{l}\}_{l=1}^{r}$ for $r\ll n$, with inducing locations ${\bf{z}}=\{{\bf{z}}_{l}\}_{l=1}^{r}$ such that ${\bf{u}}_{l}=f({\bf{z}}_{l})$. Given $\bf{u}$}, RCSVGP constructs an approximate posterior $q({\bf{u}})={\mathcal{N}}({\boldsymbol{\mu}}_{\bf{u}},{\boldsymbol{\Sigma}}_{\bf{u}})$ with variational parameters ${\boldsymbol{\mu}}_{\bf{u}}$ and ${\boldsymbol{\Sigma}}_{\bf{u}}$, obtained by maximizing the evidence lower bound (ELBO) criterion:
@@
Computation-aware GP. [wenger2022posterior, wenger2024computation] proposed computation-aware GP (CaGP) to address the SVGP overconfidence. CaGP mitigates this issue by ensuring its posterior variance exceeds the vanilla GP variance. This corrected variance is obtained by adding computational uncertainty, associated with the approximation error in the representer weights ${\bf{v}}_{\ast}=({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})^{-1}({\bf{y}}-{\bf{m}})$. To obtain computational uncertainty, CaGP places a prior $p({\bf{v}}_{\ast})={\mathcal{N}}({\bf{v}}_{0},{\boldsymbol{\Sigma}}_{0})$ and updates its posterior distribution $\mathcal{N}({\bf{v}}_{i},{\boldsymbol{\Sigma}}_{i})$ using a probabilistic linear solver based on linear Gaussian identities:
@@
where ${\bf{C}}_{i}={\bf{S}}_{i}({\bf{S}}_{i}^{\top}\hat{{\bf{K}}}^{-1}{\bf{S}}_{i}){\bf{S}}_{i}^{\top}\approx{\bf{K}}^{-1}$ is a low-rank approximation and ${\bf{S}}_{i}\in^{n\times i}$ is the action matrix associated with the chosen approximation technique. In the case of inducing points, ${\bf{S}}_{i}=k({\bf{X}},{\bf{z}}_{i})$. Given a new data point $\bf{x}$}, the predictive distribution of $f({\bf{x}})$ is Gaussian, characterized by the mean $\mu_{i}({\bf{x}})$ and variance $k_{i}({\bf{x}},{\bf{x}})$, given by:
@@ 3.1 Robust Computation-aware Inference
Inspired by CaGP, we apply a probabilistic treatment to the representer weights $\hat{{\bf{v}}}=({\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}\!-\!{\bf{m}}_{\bf{w}})$ of RCGP in Equation˜1. This enables us to quantify uncertainty in approximating $\hat{{\bf{v}}}$—an aspect most approximate GPs ignore. RCaGP later incorporates this uncertainty into its predictive variance. We begin by placing a prior on $\hat{{\bf{v}}}$, i.e., $\hat{{\bf{v}}}\sim{\mathcal{N}}(\tilde{{\bf{v}}}_{0}\!=\!\mathbf{0},\,\tilde{{\boldsymbol{\Sigma}}}_{0}\!=\!\tilde{{\bf{K}}}^{-1})$, where $\tilde{{\bf{K}}}={\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}}$. We then update this belief via the linear Gaussian identity [wenger2022posterior], yielding a Gaussian posterior with mean $\tilde{{\bf{v}}}_{i}$ and variance $\tilde{{\boldsymbol{\Sigma}}}_{i}$:
@@
where $\tilde{{\bf{C}}}_{i}\!=\!{\bf{S}}_{i}({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}$ and ${\bf{S}}_{i}\!\in\!^{n\times i}$ are low-rank approximations of $\tilde{{\bf{K}}}^{-1}$ and the actions, respectively. The updates can be done sequentially [wenger2022posterior, Algorithm 1] or in batches [wenger2024computation, Algorithm S2].
Recent work generalized the action matrix ${\bf{S}}_{i}\in^{n\times i}$ by treating it as a variable that can be optimized during model selection [wenger2024computation]. Given a new data point ${\bf{x}}\in{\mathcal{X}}$, the predictive distribution of $f({\bf{x}})$ follows a normal distribution with mean $\hat{\mu}_{i}({\bf{x}})$ and variance $\hat{k}_{i}({\bf{x}},{\bf{x}})$:
@@
In Equation˜11, RCaGP incorporates the computational uncertainty ${\bf{k}}_{\bf{x}}^{\top}\,\tilde{{\boldsymbol{\Sigma}}}_{i}\,{\bf{k}}_{\bf{x}}$, corresponding to the variance of the representer weights $\tilde{{\boldsymbol{\Sigma}}}_{i}$, resulting in a combined uncertainty $\hat{k}_{i}({\bf{x}},{\bf{x}})$. This computational uncertainty guarantees that the predictive posterior variance of RCaGP is larger than RCGP, preventing RCaGP from overconfident predictions. Such behavior is referred to as a conservative uncertainty estimate. The computational complexity of RCaGP can be found in Appendix˜B.
The robustness of RCaGP hinges on the weight function $w$, which we choose to incorporate into the representer weights $\tilde{{\bf{v}}}$. To that end, we follow [altamirano2023robust] and let:
@@
[…]
@@ 3.2 Model hyperparameters optimization
We use the evidence lower bound (ELBO) as a loss function to optimize the kernel hyperparameters ${\boldsymbol{\theta}}\!\in\!^{p}$. Following [wenger2024computation], this enables RCaGP to scale to large datasets while avoiding overconfidence. The variational family is defined using the RCaGP posterior $q_{i}({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})$, i.e., $\{q_{i}({\bf{f}})={\mathcal{N}}(\hat{\mu}_{i}({\bf{X}}),\hat{k}_{i}({\bf{X}},\!{\bf{X}}))|{\bf{S}}_{i}\!\in\!^{n\times i}\}$, parameterized by action ${\bf{S}}_{i}$. We replace the evidence term $\log p({\bf{y}}|{\boldsymbol{\theta}})$ with the robust loss $\log\Psi^{w}({\bf{y}},\!{\bf{f}})$ from [altamirano2023robust]. We then formulate ELBO as
@@
This loss learns the hyperparameters as if maximizing ${\mathbb{E}}_{p({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ while minimizing the computational uncertainty associated with the approximation error. The derivation of the expected loss ${\mathbb{E}}_{q({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ follows [altamirano2023robust], while the derivation of the KL term ${\mathrm{KL}}[q({\bf{f}})\|p({\bf{f}})]$ draws an analogy to CaGP. For the closed form and detailed derivation of ELBO, we refer the readers to Proposition˜C.1. We obtain the optimal hyperparameters $\boldsymbol{\theta}$$^{\ast}$ by minimizing the negative ELBO.